Hi,
I have an algebra question that states the following: (not sure where to go with this)
If h = -16t^2 + 48t represents the height of a rocket, in feet, t seconds after it was fired, when will the rocket hit the ground?
It does give a hint: "The rocket is on the ground when h = 0"
So it states that the rocket will hit the ground after _______ second(s).

Answer by mananth(12270)

At t=0, h=0, it is on ground before firing
h=-16t^2+48
h =0 when it hits the ground after firing
-16t^2+48=0
16t(-t+3)=0
/16
t(-t+3)=0
t= 0 OR t=3
It will hit after 3 seconds